i am working on my math homework and i came to this word problem and i tried it but i still couldn't figure out what the answer was nor could my boyfriend. please help me. two numbers total 57 and have a difference of 9.
Answer by nerdybill(6958)

two numbers total 57 and have a difference of 9
Let x,y be the two numbers
x+y = 57
x-y = 9
.
Add both equations together:
x+y = 57
x-y = 9
------------
2x = 66
x = 33
.
substitute above into:
x+y = 57
33+y = 57
y = 24
.
solution: 24 and 33